OpenStudy (anonymous):

What is the value of 3 • (–8) – 2 • (–5)?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@Tyler.Evan2

OpenStudy (anonymous):

ok, ummmm

OpenStudy (anonymous):

-14 i belive.......

OpenStudy (anonymous):

ru sure

OpenStudy (anonymous):

because 3 x -8 = -24, and -2 x -5= -10 so -24 - (-10)= -14........ so yah im sure

OpenStudy (anonymous):

im sure

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

put it in the calculator but if not its -14